# RadFlag: A Black-Box Hallucination Detection Method for Medical Vision Language Models

### v1

@@ Appendix D Alpha Comparison for Sentence-Level Flagging
Table 3: Comparison of Model Performance
@@ Appendix E Category-Level Breakdowns
Table 4: Summary of Medversa Findings at $\lambda=6$ ($\alpha=0.05$)
@@
Table 5: Summary of RaDialog Findings at $\lambda=4$ ($\alpha=0.05$)
@@ Appendix F Effect of Removing Flagged Sentences on Report-level Metrics
Table 6 shows that removing flagged sentences from reports does not seem to change overall report-level metric performance. The * indicates adjusted report-level metrics from removing empty reports that inflate RadCliQ scores.
Table 6: Performance Metrics (Findings Level) for Medversa and RaDialog
Model Method Avg. RadCliQ Avg. RadGraph Entity Precision Avg. RadGraph Entity Recall Avg. RadGraph Relation Precision Avg. RadGraph Relation Recall

### v2

@@ Appendix D Alpha Comparison for Sentence-Level Flagging
Table A2: Comparison of Model Performance
@@ Appendix E Category-Level Breakdowns
Table A3: Summary of Medversa Findings at $\lambda=6$ ($\alpha=0.05$)
@@
Table A4: Summary of RaDialog Findings at $\lambda=4$ ($\alpha=0.05$)
@@ Appendix F Effect of Removing Flagged Sentences on Report-level Metrics
Table A5 shows that removing flagged sentences from reports does not seem to change overall report-level metric performance. The * indicates adjusted report-level metrics from removing empty reports (reports with all sentences flagged) that inflate RadCliQ-v1 scores.
Table A5: Performance Metrics (Findings Level) for Medversa and RaDialog
Model Method Avg. RadCliQ-v1 Avg. RadGraph Entity Precision Avg. RadGraph Entity Recall Avg. RadGraph Relation Precision Avg. RadGraph Relation Recall
@@
Appendix G Entropy Baselines
To provide a comparison point for our black-box sampling method, we tested two baseline methods used by Manakul et al. (2023) that require access to output probabilities: average negative log probability and average entropy. In an LLM’s response, $p_{ij}$ represents the probability of generating the j-th token in the i-th sentence, where J indicates the total token count per sentence. Average negative log probability is calculated as:
$\text{Avg}(-\log p)=-\frac{1}{J}\sum_{j}\log p_{ij}$
Entropy is calculated according to the equation $\mathcal{H}_{ij}=-\sum_{\tilde{w}\in\mathcal{W}}p_{ij}(\tilde{w})\log p_{ij}(\tilde{w})$. Average entropy is thus calculated as:
$\text{Avg}(\mathcal{H})=\frac{1}{J}\sum_{j}\mathcal{H}_{ij}$
We ran experiments using a new set of MedVersa generations (temp = 0.1, excluding 11 datapoints where generation failed) on 197 MIMIC-CXR test datapoints, flagging the reports that scored highest on these metrics. Despite requiring access to output probabilities, these baseline methods performed poorly, so hallucinations appear at similar rates in ”accepted” and ”flagged” sentences.
Table A6: Comparison of different ranking methods and their hallucination detection performance.
Ranking Method Label n Mean True Hallucinations Median True Hallucinations
Average Entropy Accepted 187 2.63 2.0
Flagged 10 3.50 2.5
Accepted 174 2.59 2.0
Flagged 23 3.30 3.0
Accepted 143 2.46 2.0
Flagged 54 3.20 3.0
Average Negative Log Probability Accepted 187 2.70 2.0
Flagged 10 2.10 1.50
Accepted 174 2.59 2.0
Flagged 23 3.30 3.0
Accepted 143 2.46 2.0
Flagged 54 3.22 3.0